# Holistic Optimization Framework for FPGA Accelerators

### v3

@@ VI-B Comparison
RTL Simulation  Table II presents the RTL simulation results for Prometheus, Sisyphus [60], AutoDSE [67], ScaleHLS [79] (S-HLS), and Allo [14]. The last lines show the average and geometric mean performance improvement (PI) of Prometheus across evaluated kernels. Results marked with * are from the Vitis HLS report, showing minimum latency as an optimistic estimate. The RTL simulation for 3mm with Allo was incomplete after two days.
@@
Prometheus consistently achieves superior QoR across all evaluated kernels compared to other frameworks. While Sisyphus [60], designed primarily for computation-bound kernels, demonstrates competitive results for these kernels, it shows a weakness for 2mm and 3mm. This disparity is due to Sisyphus lacking concurrent execution capabilities for independent matrix multiplications. The performance advantage of Prometheus stems from both concurrent task execution and efficient overlapping of computation and communication.
Kernel Ours Sisyphus S-HLS Allo AutoDSE
2mm 308.38 195.09 37.13 46.58 0.41*
3mm 368.36 178.97 43.04 60.40* 1.74*
Atax 3.56 2.32 1.58 1.96 1.97*
Bicg 15.41 2.32 1.70 14.17 0.99*
Gemm 419.14 227.09 40.53 37.50 110.81*
Gesummv 10.21 2.28 1.78 8.85 1.98*
Mvt 14.65 5.54 7.39 8.77 7.80*
Symm 212.20 200.30 0.06 16.72 14.68*
Syr2k 133.77 75.01 0.04 20.59 12.35*
Syrk 316.25 211.00 0.54 41.11 23.16*
Trmm 193.47 166.72 0.07 5.10 0.02*
PI (Avg) 2.39x 927.20x 8.58x 973.14
PI (gmean) 2.03x 48.03x 4.92x 25.82
TABLE II: Comparison of the throughput (GF/s) of various optimization framework with RTL Simulation
@@
On Board Evaluation  Table III presents the results for the on-board evaluation. The column $T(ms)$ indicates the kernel execution time in milliseconds, GF/s represents the throughput in Giga Floating Operations per second, and the resource usage is detailed in the thousands for both LUTs (L) and FFs. URAM utilization is excluded as no kernels use it. Column F (MHz) shows the frequency achieved by each design, with a target frequency of 220 MHz for all designs.
@@
TABLE III: On-board evaluation comparison
@@
Table IV shows the parameters found by the NLP. We use the same name of the iterator as Polybench 4.2.1 [56] code. $Si$ represent the statement in position $i$ in the code. The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task is precise.
@@
TABLE IV: Fusion, loop order and data-tile size found by the NLP for the kernel evaluated on Table III for 1 SLR
@@
VII Related Work
@@
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks [34], Dennis Dataflow [22], synchronous dataflow languages [42, 11], and for FPGA applications [53, 12, 3, 25, 10, 5, 75, 14]. DaCe [10] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Flower [5] automates FPGA dataflow development but limits parallelism. Frameworks like [14, 75], built on HeteroCL [38], optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays [71, 20, 33, 39, 9] offer efficient computation for specific patterns but lack generalization. Application-specific frameworks [16, 29, 51, 21, 63, 23, 36, 6] demonstrate dataflow advantages but do not generalize across domains.
@@
Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion [58, 59, 60], bottleneck DSE [67], and GNN-based latency and resource estimation [66, 64, 65, 8, 68, 73, 74]. However, these approaches lack effective integration of dataflow optimization techniques.
@@
Code Transformation  Code transformation has been explored for CPUs [13, 55, 37, 7], GPUs [70], and FPGAs [57, 84, 83, 44, 48, 47, 46, 17]. Pluto [13] minimizes communication and improves locality but can limit parallelism. FPGA-specific adaptations [57] leverage FIFOs for overlapping communication and computation but are restricted in parallelism. Recent works [84, 83] selectively use Pluto for latency minimization while avoiding non-HLS-friendly code. While [44, 48, 47, 46, 17] focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives. The [79, 38, 31, 14, 78] compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties. The paper [60] described in Section II generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution. Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
@@
Tiling and Padding  Tiling is essential for balancing computation and communication. While prior works [49, 45] use cost models to minimize communication, our approach extends this to reduce overall latency. Techniques like NLP-based tiling [45] focus on CPUs, while Wedler [61] optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts. Padding is well-studied for reducing cache misses [30, 52] and improving memory transfers [72], but its use for varying unroll factors on FPGAs remains underexplored.
VIII Conclusion
In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework. By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to maximize throughput within stringent resource constraints. Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.

### v4

@@ 6.3. Comparison
Table 5 presents the RTL simulation results for Prometheus, Sisyphus (Pouget et al., 2025b), AutoDSE (Sohrabizadeh et al., 2021), ScaleHLS (Ye et al., 2022), and Allo (Chen et al., 2024). The last lines show the average and geometric mean performance improvement (PI) of Prometheus across evaluated kernels.
Kernel Ours Sisyphus ScaleHLS Allo AutoDSE Stream-HLS
2mm 308.38 195.09 37.13 46.58 0.41* 150.27
3mm 368.36 178.97 43.04 60.40* 1.74* 174.75
Atax 3.56 2.32 1.58 1.96 1.97* 1.71
Bicg 15.41 2.32 1.70 14.17 0.99* 1.73
Gemm 419.14 227.09 40.53 37.50 110.81* 203.48
Gesummv 10.21 2.28 1.78 8.85 1.98* 1.72
Mvt 14.65 5.54 7.39 8.77 7.80* 13.31
Symm 212.20 200.30 0.06 16.72 14.68* N/A
Syr2k 267.31 149.89 0.08 41.15 12.35* N/A
Syrk 158.25 105.59 0.27 20.57 23.16* N/A
Trmm 193.47 166.72 0.07 5.10 0.02* N/A
PI (Avg) 2.39x 927.20x 8.58x 973.14 3.46x
PI (gmean) 2.03x 48.03x 4.92x 25.82 2.71x
Table 5. Throughput Comparison (in GF/s) of PolyBench Kernels Across Frameworks Using RTL Simulation
Results marked with * are from the Vitis HLS report, showing minimum latency as an optimistic estimate. The RTL simulation for 3mm with Allo was incomplete after two days. The N/A values in the table for Scale-HLS correspond to kernels that have at least one loop with a non-constant trip count. Stream-HLS does not handle these cases, so we cannot make a comparison.
@@
Prometheus consistently achieves superior QoR across all evaluated kernels compared to other frameworks. While Sisyphus (Pouget et al., 2025b), designed primarily for computation-bound kernels, demonstrates competitive results for these kernels, it shows a weakness for 2mm and 3mm. This disparity is due to Sisyphus lacking concurrent execution capabilities for independent matrix multiplications. The performance advantage of Prometheus stems from both concurrent task execution and efficient overlapping of computation and communication.
@@
On Board Evaluation  Table 6 presents the results for the on-board evaluation. The column $T(ms)$ indicates the kernel execution time in milliseconds, GF/s represents the throughput in Giga Floating Operations per second, and the resource usage is detailed in the thousands for both LUTs (L) and FFs. URAM utilization is excluded as no kernels use it. Column F (MHz) shows the frequency achieved by each design, with a target frequency of 220 MHz for all designs.
@@
Table 6. On-Board Evaluation of Performance and Resource Utilization Across Frameworks and SLR Configurations
@@
Table 7 shows the parameters found by the NLP. We use the same name of the iterator as Polybench 4.2.1 (Pouchet and Yuki, 2024) code. $Si$ represent the statement in position $i$ in the code. The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task (defined in the second column) is precise.
@@
Table 7. Fusion, loop order and data-tile size found by the NLP for the kernel evaluated on Table 6 for 1 SLR
@@
6.4. Scalability
Our NLP solver includes the option to set a timeout, returning the best design found so far without guaranteeing optimality. This feature enables efficient exploration of the solution space while ensuring adherence to time constraints when necessary.
When analyzing the time required to solve the Non-Linear Programming (NLP) problem in Sisyphus, we observe that our framework, Prometheus, achieves similar solution times for most benchmarks. However, there is a notable exception: the 3mm benchmark, which times out after 4 hours in Sisyphus, while Prometheus successfully finds a solution.
This efficiency stems from Prometheus’ ability to explore a larger optimization space while still maintaining fast solution times. The key reason is that all optimization techniques are seamlessly integrated within the design space. This structured integration ensures that when a decision is made for one optimization parameter, it naturally constrains the possible choices for others, thereby reducing the search complexity. As a result, even with a broader search space, Prometheus efficiently converges to a solution in 21.37 seconds.
A detailed comparison of solution times can be found in Table 8, which highlights the impact of our approach in managing the design space effectively.
Benchmark Sisyphus Prometheus
2mm 22.37 15.98
3mm 14400.08 21.37
Atax 3.18 7.03
Bicg 3.17 5.09
Gemm 2.19 4.31
Gesummv 1.57 8.07
Mvt 1.71 1.26
Symm 7.13 11.46
Syr2k 5.29 6.07
Syrk 2.56 3.94
Trmm 4.34 5.96
Average 1313.96 8.23
Geo Mean 7.95 6.50
Table 8. Time Required (in seconds) for the NLP Solver to Find an Optimal Solution: Sisyphus vs. Prometheus Across Benchmarks (Timeout Set at 14,400s)
7. Related Work
@@
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks (Kahn, 1974), Dennis Dataflow (Dennis, 1974), synchronous dataflow languages (Lee and Messerschmitt, 1987; Benveniste et al., 1994), and for FPGA applications (Peverelli et al., 2018; Bhattacharyya et al., 2009; Abid et al., 2013; Grigoraş et al., 2013; Ben-Nun et al., 2019; Amiri et al., 2021; Xiang et al., 2022; Chen et al., 2024). DaCe (Ben-Nun et al., 2019) introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Stream-HLS (Basalama and Cong, 2025) automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory. Additionally, it offers very limited opportunities for parallelism. Flower (Amiri et al., 2021) automates FPGA dataflow development but limits parallelism. Frameworks like (Chen et al., 2024; Xiang et al., 2022), built on HeteroCL (Lai et al., 2019), optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays (Wang et al., 2021; Cong and Wang, 2018; Jia et al., 2020; Lai et al., 2020; Basalama et al., 2023) offer efficient computation for specific patterns but lack generalization. Application-specific frameworks (Chi et al., 2018; Hegarty et al., 2014; Natale et al., 2016; de Fine Licht et al., 2021; Singh et al., 2020; Denzler et al., 2023; Korol et al., 2022; Bacis et al., 2017) demonstrate dataflow advantages but do not generalize across domains. RapidStream-TAPA (Guo et al., 2023a, b) enhances the performance of dataflow designs and automates SLR placement. However, it requires an optimized kernel with a dataflow structure as input.
@@
Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion (Pouget et al., 2024, 2025a, 2025b), bottleneck DSE (Sohrabizadeh et al., 2021), and GNN-based latency and resource estimation (Sohrabizadeh et al., 2023, 2022a, 2022b; Bai et al., 2022; Ustun et al., 2020; Wu et al., 2022a, b). However, these approaches lack effective integration of dataflow optimization techniques.
@@
Code Transformation  Code transformation has been explored for CPUs (Bondhugula et al., 2008; Pouchet et al., 2011; Kruse et al., 2020; Baghdadi et al., 2019), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018). Pluto (Bondhugula et al., 2008) minimizes communication and improves locality but can limit parallelism. FPGA-specific adaptations (Pouchet et al., 2013b) leverage FIFOs for overlapping communication and computation but are restricted in parallelism. Recent works (Zhao et al., 2022; Zhao and Cheng, 2021) selectively use Pluto for latency minimization while avoiding non-HLS-friendly code. While (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives. The (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021; Chen et al., 2024; Ye et al., 2024) compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties. The paper (Pouget et al., 2025b) described in Section 2 generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution. Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
@@
Tiling and Padding  Tiling is essential for balancing computation and communication. While prior works (Liu et al., 2017a; Li et al., 2021) use cost models to minimize communication, our approach extends this to reduce overall latency. Techniques like NLP-based tiling (Li et al., 2021) focus on CPUs, while Wedler (Shi et al., 2023) optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts. Padding is well-studied for reducing cache misses (Hong et al., 2016; Panda et al., 1999) and improving memory transfers (Wang et al., 2012), but its use for varying unroll factors on FPGAs remains underexplored.
8. Conclusion
In this work, we introduced Prometheus, a holistic optimization framework that unifies loop transformations, task concurrency, computation-communication overlap, and hardware-aware scheduling for FPGA accelerators. By formulating the optimization process as a Non-Linear Programming (NLP) problem, Prometheus enables a structured exploration of the vast design space while considering hardware constraints, memory bandwidth, and task parallelism.
Our framework addresses key limitations in existing methodologies, which often optimize only isolated aspects of FPGA acceleration. Prometheus surpasses these approaches by integrating SLR-aware scheduling, dynamic memory management, and hybrid execution models that effectively balance dataflow streaming and shared buffering. This enables more efficient utilization of FPGA resources, leading to significant improvements in latency, throughput, and scalability.
Through extensive performance evaluations on computation-bound kernels, we demonstrated that Prometheus outperforms state-of-the-art frameworks.
Furthermore, Prometheus’ automatic design space exploration significantly reduces the manual effort required in FPGA development. By automatically generating HLS-C++ code, OpenCL host code, and FPGA bitstreams, the framework streamlines the deployment process, making FPGA acceleration more accessible to a broader range of applications.
In summary, Prometheus provides an innovative and effective approach to FPGA optimization, delivering high-performance solutions while minimizing the complexity of hardware design.
Acknowledgments
This work was supported by the NSF award #CCF-2211557. It is also supported by CDSC industrial partners and the AMD/HACC Program.